Belton Company Currently sells its products for $25 per unit. Management is contemplating a 20% increase in the sale price for the next year. Variable costs are currently 30% of sales revenue and are not expected to change next year. Fixed expenses are 150,000 per year. If fixed costs increase 10% next year, and the new sale price per unit goes into effect, how many units will need to be sold to breakeven?
Answer #1

New fixed cost = 150,000 + 10% = 165,000

New selling price per unit = 25 + 20% = 30

New variable cost = 30 * 30% = 9

Breakeven point = Fixed cost / Contribution margin per unit

= 165,000/(30-9)

= 7857 units